Write an equation in standard form of the line passing through (2,-5) and (-4,7) Thanks. Please answer ASAP for I have a test on monday.
I have tried doing the first y minus second y over first x minus second x to find the slope. I got -2. Then I plugged it in to the slope intercept form. Y =mx +b I plugged in one ordered pair and got the y intercept is equal to -1. I got -2x + y = -1 but the textbook said that the answer was 2x +y =-1

Answer by mananth(16946)

Write an equation in standard form of the line passing through (2,-5) and (-4,7) Thanks. Please answer ASAP for I have a test on monday.
(2,-5) and (-4,7)
The slope of line passing through (x1,y1),(x2,y2) = (y2-y1)/(x2-x1)
m=7-(-5)/(-4-2)
m=12/-6
m=-2
...
y= mx +b .
take any point above
(2,-5)
plug the point and slope to find b
-5=-2*2+b
-5+4=b
b=-1
so the equation is
y=-2x-1